# On the Foundation of Distributionally Robust Reinforcement Learning

### v3

@@ 2.2.2 S-Rectangular Set of Adversary’s Policies
Consider an inventory control problem where the next day’s inventory level is
$I_{t+1}=(I_{t}+a_{t}+D_{t})_{+}$
where $\left\{{D_{t}:t\geq 0}\right\}$ i.i.d. is the demand process and $a_{t}$ is the ordered inventory. In this modeling environment, it is natural to assume that at each time $t$, the adversary can only change the law of $D_{t}$ dependent on $I_{t}$ but not dependent on the controller’s action $a_{t}$. This leads to an S-rectangular set of adversarial decision rules. However, if we instead assume SA-rectangularity, then the adversary must be able to freely choose different laws for $D_{t}$ for different controller’s action $a$; say if $a_{t}$ is large, then the adversary chooses the demand to be small, and if $a_{t}$ is small, chooses the demand to be large. Through this example, one can see that the S-rectangularity allows the modeler to limit the power of the adversary so that it cannot use the realization of the controller’s action.
Literature explores risk-averse distributionally robust control of inventory models, employing a wide array of adversaries and establishing properties of the optimal controller; for instance Shapiro and Xin (2020) and Shapiro (2022).
From these definitions, it is evident that an SA-rectangular set of adversarial policies is S-rectangular with
$\mathcal{P}_{s}:=\left\{{p_{s}=\left\{{p_{s,a}:a\in A}\right\}:p_{s,a}\in\mathcal{P}_{s,a}}\right\}=\prod_{a\in A}\mathcal{P}_{s,a}$ (2.8)
@@ 2.2.3 General Rectangular Set of Adversary’s Policies
For completeness, we also introduce the general (a.k.a. non-rectangular in the literature, see Wiesemann et al. (2013)) sets of adversarial decision rules. In this case, choosing adversarial decision rule $\kappa_{t}(g_{t-1},s,a)=p_{s,a}$ for one pair of $(s,a)\in S\times A$ could affect the possible choice of $\kappa_{t}(g_{t-1},s^{\prime},a^{\prime})$ for any other $(s^{\prime},a^{\prime})\in S\times A$.
@@
$\mathrm{K}_{\mathrm{H}}^{\mathrm{N}}(t):=\left\{{\kappa_{t}:\kappa_{t}(g_{t-1},\cdot,\cdot)\in\mathcal{P},\forall g_{t-1}\in\mathbf{G}_{t-1}}\right\}.$ (2.9)
The general history-dependent/Markov/Markov time-homogeneous adversarial policy classes $\mathrm{K}_{\mathrm{H}}^{\mathrm{N}}/\mathrm{K}_{\mathrm{M}}^{\mathrm{N}}/\mathrm{K}_{\mathrm{S}}^{\mathrm{N}}$ can be defined analogously. We omit the presentation.
@@ 2.2.4 Adversary’s Policy: Summary of Notations
Adversarial Action $\kappa_{t}(g_{t-1},s,a)=p_{s,a}\in\mathcal{P}_{s,a}$ $\kappa_{t}(g_{t-1},s,\cdot)=p_{s}\in\mathcal{P}_{s}$ $\kappa_{t}(g_{t-1},\cdot,\cdot)=p\in\mathcal{P}$
@@
It’s crucial to discern from Table 5 that the concept of rectangularity is closely tied to the extent of historical information available to the adversary before making a decision. To illustrate, consider the following scenarios: An SA-rectangular adversary at time $t$ possesses visibility into $g_{t}=(g_{t-1},s_{t},a_{t})$ and hence can be seen as making a choice from the action set $p_{s_{t},a_{t}}\in\mathcal{P}_{s_{t},a_{t}}$. In contrast, an S-rectangular adversary sees only $(g_{t-1},s_{t})$ and must determine the distribution of the next state for each possible action $a_{t}$. Consequently, it selects an action $p_{s_{t}}\in\mathcal{P}_{s_{t}}$. It is possible to further reduce the amount of historical information accessible to the adversary at the time of decision making, leading to the definition of even more general adversarial policy classes. However, such settings typically devoid the DPP and are therefore beyond the scope of this paper’s concerns. We remark that this perspective of information availability is more explicitly represented by the multistage stochastic program formulation in Section 6.1.
@@ 2.3 The Max-Min Control Problem
With the careful establishment of the controller’s and adversary’s policy classes in this section, we lay the foundation for the subsequent definition of the max-min control value of a robust MDP. The formulation of the value function allows us to formalize robust policy learning and decision-making using the robust MDP framework.
We start with introducing some notations. For function $f$ on the measurable space $(E,\mathcal{E})$ and measure $\nu\in\mathcal{P}(\mathcal{E})$, we define the integral
$\nu[f]:=\sum_{e\in E}f(e)\nu(\left\{{e}\right\}),$ (2.13)
which is equivalent to the expectation $E_{Z\sim\nu}[f(Z)]$. Through an abuse of notation, we write $\nu[e]=\nu(\left\{{e}\right\})$ when $e\in E$, equivalent to the probability of event $e$ under measure $\nu$. Moreover, for $d\in\mathcal{P}(\mathcal{A})$ and $p_{s}=\left\{{p_{s,a}:a\in A}\right\}$, we define the measure $d\otimes p_{s}\in\mathcal{P}(\mathcal{S}\times\mathcal{A})$ by
$d\otimes p_{s}[f]=\sum_{a\in A}d[a]\sum_{s^{\prime}\in S}p_{s,a}[s^{\prime}]f(s,a)$ (2.14)
for all $f:S\times A\rightarrow\mathbb{R}$. Of course, when $u:S\rightarrow\mathbb{R}$, we have
$d\otimes p_{s}[u]=d\otimes p_{s}[u\mathds{1}_{a\in A}]=\sum_{a\in A}d[a]p_{s,a}[u].$
With these notations in place, we are now equipped to articulate the collective effect of a pair $(\pi,\kappa)$, comprising the controller’s and adversary’s policies, from the point of view of the probabilistic characteristics of the state-action sequence. This effect is manifested by uniquely defining the distribution of the process $\left\{{(X_{t},A_{t}):t\geq 0}\right\}$.
@@
For any $\pi\in\Pi_{\mathrm{H}}$, $\kappa\in\mathrm{K}_{\mathrm{H}}$, and initial distribution $\mu\in\mathcal{P}(\mathcal{S})$, define $P^{\pi,\kappa}_{\mu}$ to be the unique probability measure on $(\Omega,\mathcal{F})$ such that
$P^{\pi,\kappa}_{\mu}(G_{t}(\omega)=g_{t})=\mu(s_{0})d_{0}(s_{0})[a_{0}]\kappa_{0}(s_{0},a_{0})[s_{1}]d_{1}(s_{0},a_{0},s_{1})[a_{1}]\kappa_{1}(s_{0},a_{0},s_{1},a_{1})[s_{2}]\dots d_{t}(h_{t})[a_{t}]$ (2.15)
@@
We remark that (2.15) defines a measure on the cylinder sets on $\Omega$, hence uniquely extends to a measure on $(\Omega,\mathcal{F})$. Also, note that (2.15) determines the finite dimensional distribution of the stochastic process $\left\{{(X_{t},A_{t}):t\geq 0}\right\}$.
Now, we are ready to define the max-min control value of a robust MDP problem. Let $r:S\times A\rightarrow[0,1]$ be a bounded non-negative reward function, and $\gamma\in(0,1)$ be the discount factor.
@@
For any subsets $\Pi\subset\Pi_{\mathrm{H}}$ and $\mathrm{K}\subset\mathrm{K}_{\mathrm{H}}$, we define the max-min control optimal value associated with the triple $(\mu,\Pi,\mathrm{K})$ by
$v^{*}(\mu,\Pi,\mathrm{K}):=\sup_{\pi\in\Pi}\inf_{\kappa\in\mathrm{K}}v(\mu,\pi,\kappa).$ (2.16)
To familiarize the reader with this definition in conjunction with the policy classes of the controller and the adversary, we now introduce the following set of elementary properties of the max-min control optimal value.
@@
Recall the definitions of the controller’s and adversary’s policy classes. Then,
$v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{H}},\mathrm{K})\geq v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{M}},\mathrm{K})\geq v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{S}},\mathrm{K})$
for $\mathrm{K}=\mathrm{K}_{\mathrm{H}}^{\mathrm{S}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$ and
$v^{*}(\mu,\Pi,\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})\leq v^{*}(\mu,\Pi,\mathrm{K}_{\mathrm{M}}^{\mathrm{S}})\leq v^{*}(\mu,\Pi,\mathrm{K}_{\mathrm{S}}^{\mathrm{S}})$
for $\Pi=\Pi^{\mathrm{C}}_{\mathrm{H}},\Pi^{\mathrm{C}}_{\mathrm{M}},\Pi^{\mathrm{C}}_{\mathrm{S}}$ as defined in (2.1).
This lemma follows from the inclusion relationships of the policy classes, see Appendix B.1 for a proof. Note that since a SA-rectangular set of adversarial policies is S-rectangular as explained in (2.8), Lemma 1 is also true if we replace $\mathrm{K}_{\mathrm{H}}^{\mathrm{S}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$ with $\mathrm{K}_{\mathrm{H}}^{\mathrm{SA}},\mathrm{K}_{\mathrm{M}}^{\mathrm{SA}},\mathrm{K}_{\mathrm{S}}^{\mathrm{SA}}$.
@@ 3 Dynamic Programming Principles
As previously highlighted, the motivation behind defining rectangular sets of adversarial policies is to ensure that the max-min control problem adheres to a DPP. This is manifested in the form of the value function being the solution to the DR Bellman equation, a concept that we shall define next. We have noted that it is in general impossible to formulate a DPP for the general rectangular adversary. So, in the following sections, we will restrict our attention to the S-rectangular case.
Definition 3 (DR Bellman Equation).
Given the constraint set of action distributions $\mathcal{Q}\subset\mathcal{P}(\mathcal{A})$ and S-rectangular adversarial action sets $\mathcal{P}_{s}\subset\left\{{A\rightarrow\mathcal{P}(\mathcal{S})}\right\}$, $s\in S$, define the distributionally robust Bellman equation as
$u(s)=\sup_{d\in\mathcal{Q}}\inf_{p_{s}\in\mathcal{P}_{s}}d[r(s,\cdot)]+\gamma d\otimes p_{s}[u],\quad s\in S.$ (3.1)
Moreover, we interchange the infimum and supremum to introduce the following inf-sup equation. This equation plays a crucial role in deducing sufficient conditions for the optimal value $v^{*}$ to satisfy a DPP:
$u(s)=\inf_{p_{s}\in\mathcal{P}_{s}}\sup_{d\in\mathcal{Q}}d[r(s,\cdot)]+\gamma d\otimes p_{s}[u],\quad s\in S.$ (3.2)
In the context of finite state and action spaces DRMDPs, a bounded solution of the DR Bellman equation (3.1) always exists and is unique:
@@
This proposition follows from the well-known contraction property of the distributionally robust Bellman operator. We include a proof in Appendix B.2 so as to make the work self-contained.
Having confirmed the existence of a solution to the Bellman equation, we define the satisfaction of the DPP as the max-min optimal value being identified with that solution in the following sense.
@@
Given the action distribution sets $\mathcal{Q},\left\{{\mathcal{P}_{s}:s\in S}\right\}$, we say that the pair of controller and adversary’s policy class $(\Pi,\mathrm{K})$ satisfies the dynamic programming principle (DPP) if the solution $u^{*}$ to the DR Bellman equation (3.1) satisfies $\mu[u^{*}]=v^{*}(\mu,\Pi,\mathrm{K})$ for all $\mu\in\mathcal{P}(\mathcal{S})$.
As highlighted in the introduction, the significance of the satisfaction of a DPP in the context of DRRL lies in the fundamental role played by the Bellman equation, which serves as the primary computational tool underlying nearly all RL algorithms. In particular, there is no known computationally tractable algorithms that can achieve policy learning within the DRRL framework where the solution of the Bellman equation doesn’t correspond to the max-min optimal value.
@@
3.1 Maxmin Optimal Values and Bellman Equation: The General Case
@@
$\mu[u^{*}]$ $=v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{H}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})$
$=v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{M}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})$ $=v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{M}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}})$
$=v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{S}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})$ $=v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{S}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}})$ $=v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{S}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}})$
for all $\mu\in\mathcal{P}(\mathcal{S})$.
The proof of Theorem 1 is deferred to Appendix A.1.

### v4

@@ 2.2.2 S-Rectangular Set of Adversary’s Policies
Consider an inventory control problem where the next day’s inventory level is $I_{t+1}=(I_{t}+a_{t}+D_{t})_{+}$ where $\left\{{D_{t}:t\geq 0}\right\}$ i.i.d. is the demand process and $a_{t}$ is the ordered inventory. In this modeling environment, it is natural to assume that at each time $t$, the adversary can only change the law of $D_{t}$ dependent on $I_{t}$ but not dependent on the controller’s action $a_{t}$. This leads to an S-rectangular set of adversarial decision rules. However, if we instead assume SA-rectangularity, then the adversary must be able to freely choose different laws for $D_{t}$ for different controller’s action $a$; say if $a_{t}$ is large, then the adversary chooses the demand to be small, and if $a_{t}$ is small, chooses the demand to be large. Through this example, one can see that the S-rectangularity allows the modeler to limit the power of the adversary so that it cannot use the realization of the controller’s action.
Comparing the definition of $\mathcal{P}_{s}$ and $\mathcal{P}_{s,a}$, it is evident that an SA-rectangular set of adversarial policies is S-rectangular with
$\mathcal{P}_{s}:=\left\{{p_{s}=\left\{{p_{s,a}:a\in A}\right\}:p_{s,a}\in\mathcal{P}_{s,a}}\right\}=\bigtimes_{a\in A}\mathcal{P}_{s,a}$ (2.8)
@@ 2.2.3 General Rectangular Set of Adversary’s Policies
For completeness, we also introduce the general (a.k.a. non-rectangular in the literature, see Wiesemann et al., [79]) sets of adversarial decision rules. In this case, choosing adversarial decision rule $\kappa_{t}(\cdot|g_{t-1},s,a)=p_{s,a}$ for one pair of $(s,a)\in S\times A$ could affect the possible choice of $\kappa_{t}(\cdot|g_{t-1},s^{\prime},a^{\prime})$ for any other $(s^{\prime},a^{\prime})\in S\times A$.
@@
$\mathrm{K}_{\mathrm{H}}^{\mathrm{N}}(\mathcal{P},t):=\left\{{\kappa_{t}:\kappa_{t}(\cdot|g_{t-1},\cdot,\cdot)\in\mathcal{P},\forall g_{t-1}\in\mathbf{G}_{t-1}}\right\}.$ (2.9)
The general history-dependent/Markov/Markov time-homogeneous adversarial policy classes $\mathrm{K}_{\mathrm{H}}^{\mathrm{N}}/\mathrm{K}_{\mathrm{M}}^{\mathrm{N}}/\mathrm{K}_{\mathrm{S}}^{\mathrm{N}}$ can be defined analogously. We omit the presentation. We remark that the term “general rectangular” is used to highlight that the adversary remains rectangular in the sense that the history $g_{t-1}$ does not influence the set $\mathcal{P}$ from which the adversary selects its action.
@@ 2.2.4 Adversary’s Policy: Summary of Notations
Adversarial Action $\kappa_{t}(\cdot|g_{t-1},s,a)\in\mathcal{P}_{s,a}$ $\kappa_{t}(\cdot|g_{t-1},s,\cdot)\in\mathcal{P}_{s}$ $\kappa_{t}(\cdot|g_{t-1},\cdot,\cdot)\in\mathcal{P}$
@@
It’s crucial to realize from Table 5 that the concept of rectangularity is connected to the extent of historical information available to the adversary before making a decision. To illustrate, consider the following scenarios: An SA-rectangular adversary at time $t$ possesses visibility of $g_{t}=(g_{t-1},s_{t},a_{t})$ and hence can be seen as making a choice from the action set $p_{s_{t},a_{t}}\in\mathcal{P}_{s_{t},a_{t}}$. In contrast, an S-rectangular adversary sees only $(g_{t-1},s_{t})$ and must determine the distribution of the next state for each possible action $a_{t}$. Consequently, it selects an action $p_{s_{t}}\in\mathcal{P}_{s_{t}}$. It is possible to further reduce the amount of historical information accessible to the adversary at the time of decision making, leading to the definition of even more general adversarial policy classes. However, such settings typically lack a valid DPP and are therefore beyond the scope of this paper. We remark that this perspective of information availability is more explicitly represented by the multistage stochastic program formulation in Section 7.1.
@@ 2.3 The Max-Min Control Problem
With the careful establishment of the controller’s and adversary’s policy classes in this section, we lay the foundation for the subsequent definition of the max-min control value of an RMDP. The formulation of the value function allows us to formalize robust policy learning and decision-making using the RMDP framework.
Having rigorously defined the controller’s policy $\pi$ and adversary’s policy $\kappa$, we characterize the probability model for the state-action sequence $\left\{{(X_{t},A_{t}):t\geq 0}\right\}$ under the policy pair $(\pi,\kappa)$.
@@
For any pair of policies $\pi=(\pi_{0},\pi_{1},\dots)$ and $\kappa=(\kappa_{0},\kappa_{1},\dots)$, and initial distribution $\mu\in\mathcal{P}(\mathcal{S})$, define $P^{\pi,\kappa}_{\mu}$ to be the unique probability measure on $(\Omega,\mathcal{F})$ such that
$P^{\pi,\kappa}_{\mu}(G_{t}(\omega)=g_{t})=\mu(s_{0})\pi_{0}(a_{0}|s_{0})\kappa_{0}(s_{1}|s_{0},a_{0})\pi_{1}(a_{1}|s_{0},a_{0},s_{1})\kappa_{1}(s_{2}|s_{0},a_{0},s_{1},a_{1})\cdots\pi_{t}(a_{t}|h_{t})$ (2.13)
@@
We remark that (2.13) consistently defines the measure value on the cylinder sets on $\Omega$, hence uniquely extends to a measure on $(\Omega,\mathcal{F})$ by Carathéodory’s extension theorem. Also, note that (2.13) determines the finite dimensional distribution of the stochastic process $\left\{{(X_{t},A_{t}):t\geq 0}\right\}$.
Now, we are ready to define the max-min control value of an RMDP problem. Without loss of generality, we will assume that the reward function is non-negative and bounded by 1 throughout the paper; i.e. $r:S\times A\rightarrow[0,1]$. Let $\gamma\in(0,1)$ be the discount factor.
@@
For any pair controller and adversarial policy classes $(\Pi,\mathrm{K})$ and initial distribution $\mu$, we define the max-min control optimal value by
$v(\mu,\Pi,\mathrm{K}):=\sup_{\pi\in\Pi}\inf_{\kappa\in\mathrm{K}}v(\mu,\pi,\kappa).$ (2.14)
We introduce the following set of elementary properties of the max-min control optimal value.
@@
Recall the definitions of the controller’s and adversary’s policy classes. Then, for any controller action set $\mathcal{Q}$ and S-rectangular adversary action set $\mathcal{P}^{\mathrm{S}}$, $v(\mu,\Pi_{\mathrm{H}},\mathrm{K})\geq v(\mu,\Pi_{\mathrm{M}},\mathrm{K})\geq v(\mu,\Pi_{\mathrm{S}},\mathrm{K})$ for $\mathrm{K}=\mathrm{K}_{\mathrm{H}}^{\mathrm{S}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$ and $v(\mu,\Pi,\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})\leq v(\mu,\Pi,\mathrm{K}_{\mathrm{M}}^{\mathrm{S}})\leq v(\mu,\Pi,\mathrm{K}_{\mathrm{S}}^{\mathrm{S}})$ for $\Pi=\Pi_{\mathrm{H}},\Pi_{\mathrm{M}},\Pi_{\mathrm{S}}$ as defined in (2.1).
This lemma follows from the inclusion relationships of the policy classes, see Appendix A.1 for a proof. Note that since a SA-rectangular set of adversarial policies is S-rectangular as explained in (2.8), Lemma 1 is also true if we replace $\mathrm{K}_{\mathrm{H}}^{\mathrm{S}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$ with $\mathrm{K}_{\mathrm{H}}^{\mathrm{SA}},\mathrm{K}_{\mathrm{M}}^{\mathrm{SA}},\mathrm{K}_{\mathrm{S}}^{\mathrm{SA}}$.
@@ 3 Dynamic Programming Principles
The motivation for the use of rectangular adversarial policy classes is to ensure that the max-min control problem satisfies a DPP. Specifically, rectangularity ensures that the resulting value function satisfies the robust Bellman equation, which we define next.
Definition 3 (Robust Bellman Equation).
Given the constraint set of action distributions $\mathcal{Q}\subset\mathcal{P}(\mathcal{A})$ and S-rectangular adversarial action sets $\mathcal{P}_{s},s\in S$, define the robust Bellman equation as
$u(s)=\sup_{\phi\in\mathcal{Q}}\inf_{p_{s}\in\mathcal{P}_{s}}E_{\phi,p_{s}}\left[r(s,A_{0})+\gamma u(X_{1})\right],\quad s\in S.$ (3.1)
Here, the law of $A_{0},X_{1}$ under $E_{\phi,p_{s}}$ is defined by $P_{\phi,p_{s}}(A_{0}=a_{0},X_{1}=s_{1})=\phi(a_{0})p_{s,a_{0}}(s_{1})$.
Another equation of significant interest is the following inf-sup equation, where the supremum and infimum operations are interchanged. This equation plays a crucial role in deducing sufficient conditions for the optimal value $v^{*}$ to satisfy a DPP:
$u^{\prime}(s)=\inf_{p_{s}\in\mathcal{P}_{s}}\sup_{\phi\in\mathcal{Q}}E_{\phi,p_{s}}\left[r(s,A_{0})+\gamma u^{\prime}(X_{1})\right],\quad s\in S.$ (3.2)
In the context of finite state and action spaces RMDPs, a bounded solution of the robust Bellman equation (3.1) always exists and is unique:
@@
This proposition follows from the well-known contraction property of the robust Bellman operator (see, for example, Theorem 4 in Wiesemann et al., [79]). We include a proof in Appendix A.2 to make the work self-contained. Having confirmed the existence of $u^{*}$, we define the satisfaction of the DPP as the max-min optimal value being identifiable with $u^{*}$ in the following sense.
@@
Given the action distribution sets $\mathcal{Q}$ and $\mathcal{P}^{\mathrm{S}}=\left\{{\mathcal{P}_{s}:s\in S}\right\}$, we say that the pair of controller and adversary’s policy class $(\Pi,\mathrm{K})$ satisfies the dynamic programming principle (DPP) if the solution $u^{*}$ to the robust Bellman equation (3.1) satisfies $E_{\mu}[u^{*}(X_{0})]=v(\mu,\Pi,\mathrm{K})$ for all $\mu\in\mathcal{P}(\mathcal{S})$.
As highlighted in the introduction, the significance of the satisfaction of a DPP in the context of DRRL lies in the fundamental role played by the Bellman equation, which serves as the primary computational tool underlying nearly all RL algorithms. In particular, all existing DRRL algorithms achieve policy learning by representing the value function through fixed-point equation (3.1) or (3.2).
@@
3.1 Max-Min Optimal Values and Bellman Equation: The General Case
@@
$E_{\mu}[u^{*}(X_{0})]$ $=v(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})$
$=v(\mu,\Pi_{\mathrm{M}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})$ $=v(\mu,\Pi_{\mathrm{M}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}})$
$=v(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})$ $=v(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}})$ $=v(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}})$
for all $\mu\in\mathcal{P}(\mathcal{S})$; i.e. the DPP holds for these cases.
The proof of Theorem 1 relies on two important facts. First, the DPP always holds for controllers and adversaries with symmetric information structures. Second, $E_{\mu}[u^{*}(X_{0})]$ is always a lower bound for the optimal maxmin control values regardless of information asymmetries and constraints. With these two facts and Lemma 1, one can directly conclude Theorem 1. The detailed proof is deferred to Appendix B.1.